A normal distribution has a mean of 80 and a standard deviation of 14. Determine the value above which 80% of values will occur
Answer #1
A normal distribution has a mean of 80 and a standard deviation of 14. Determine the value above which 80% of values will occur ----------------- The z-value for 80% is inVNorm(0.8) = 0.841621... ----------------- Use the formula x =z*s + u to solve for "x": x = 0.841621*14 + 80 x = 91.783 ====================== Cheers, Stan H.
answered by: Thomas S.